The mean cost of domestic airfares in the United States rose to an all-time high of $370 per ticket. Airfares were based on the total ticket value, which consisted of the price charged by the airlines plus any additional taxes and fees. Assume domestic airfares are normally distributed with a standard deviation of $120.

a. What is the probability that a domestic airfare is $555 or more (to 4 decimals)?

b. What is the probability that a domestic airfare is $260 or less (to 4 decimals)?

c. What if the probability that a domestic airfare is between $310 and $510 (to 4 decimals)?

d. What is the cost for the 4% highest domestic airfares? (rounded to nearest dollar) $ or
Answer #1

the PDF of normal distribution is = 1/σ * √2π * e ^ -(x-u)^2/ 2σ^2
standard normal distribution is a normal distribution with a,
mean of 0,
standard deviation of 1
equation of the normal curve is ( Z )= x - u / sd ~ N(0,1)
mean ( u ) = 370
standard Deviation ( sd )= 120
------------------------------------------------------------------------------
a.
P(X < 555) = (555-370)/120
= 185/120= 1.5417
= P ( Z <1.5417) From Standard Normal Table
= 0.9384
P(X > = 555) = (1 - P(X < 555)
= 1 - 0.9384 = 0.0616
------------------------------------------------------------------------------
b.
P(X > 260) = (260-370)/120
= -110/120 = -0.9167
= P ( Z >-0.9167) From Standard Normal Table
= 0.8203,
P(X < = 260) = (1 - P(X > 260)
= 1 - 0.8203 = 0.1797
------------------------------------------------------------------------------
c.
To find P(a < = Z < = b) = F(b) - F(a)
P(X < 310) = (310-370)/120
= -60/120 = -0.5
= P ( Z <-0.5) From Standard Normal Table
= 0.3085
P(X < 510) = (510-370)/120
= 140/120 = 1.1667
= P ( Z <1.1667) From Standard Normal Table
= 0.8783
P(310 < X < 510) = 0.8783-0.3085 = 0.5698
------------------------------------------------------------------------------
d.
P ( Z > x ) = 0.04
Value of z to the cumulative probability of 0.04 from normal table is 1.750686
P( x-u / (s.d) > x - 370/120) = 0.04
That is, ( x - 370/120) = 1.750686
--> x = 1.750686 * 120+370 = 580.0823